1. A ball of mass 14 kg is rolling towards the wall at 5.00 m/s. After the impact with the wall, the ball rolls back at 4.60 m/s. Calculate the magnitude of change in momentum of the ball.

2. A typical rifle has a mass of 3.0 kg. A bullet in that rifle has a mass of 3.0 g. After the bullet is fired, it has a speed of 100 m/s. Calculate the magnitude of average force exerted by the rifle on a person's shoulder due to the rifle's recoil, if the time duration of impact with the shoulder is 1.6 s.
Answer #1

1)

Let

v1 = -5 m/s

v2 = 4.6 m/s

m = 14 kg

change in momentum = m*(v2 - v1)

= 14*(4.6 - (-5))

= 134.4 kg.m/s <<<<<<<--------------Answer

2) Impulse on the bullet

Let F_avg is the average force exerted on the bullet,

Impulse = change in momentum of the bullet

F_avg*dt = m*(v2 - v1)

F_avg = m*(v2 - v1)/dt

= 0.003*(100 - 0)/1.6

= 0.1875 kg.m/s <<<<<<<--------------Answer

so, the same amount of force is exerted by the rifle on shoulder.